13. Train tickets are sold as follows:

  • Northeast is coded A, Midwest B, South C, and West D.
  • If the train leaves between 10pm and 6am the above codes are lower case instead of upper-case letters.
  • Male passengers are coded X and female passengers are coded Y, children under the age of 14 are coded in lower case.
  • Meals are coded as follows: Salad G, Sandwich H, Pasta K. Again, children are coded in lower case.
  • First class is P, business class is Q, and economy class is R.

  

What is the code for a trip to Boston leaving at 5am for a sandwich eating 11-year-old male traveling economy class?

*

14. Three students were lined up in a row. Damon was to the left of Val but not necessarily next to them. The student wearing the blue shirt was to the right of the student wearing the white shirt. The student wearing the black shirt was to the left of Harold. Harold was to the left of Val.

   

What was the order of the students from left to right? (Names should be separated by commas - e.g. Jane, Tri, Dave)

*

15. Ming bought a large bag of apples and ate a fifth of them the first day. On the next day, she ate a quarter of the remaining apples. The following day she consumed half of the apples that were left before throwing out the last 6 apples.



  

How many apples did Ming buy?

*
Answer #1

Answer 13. As Boston is in North-east and the train leaves at 5 pm, so the code will be 'a'.

As, it is a male child with less than 11 years of age, so the code will be 'x'.

Code for sandwich for a child is 'h' and economy class is 'R'.

So, the code is 'axhR'.

Answer 14. It is given, both Harold and Damon are to the left of Val. Hence, Val is to the extreme right.

Again, it is mentioned that the student wearing the black shirt is to the left of Harold. So, as Val is to the extreme right, the only person left is Damon, so Damon is to the immediate left of Harold.

So, the order of students from left to right is : Damon, Harold, Val.

Answer 15 : It is given that Ming threw out the last apples left with her and they were 6 in number.

Let the number of apples that Ming brought be 100.

So, the number of apples she ate on the first day = 1/5\times100 = 20

So, after 1st day she had 100-20 apples = 80 apples

Number of apples she ate on 2nd day = 1/4\times80 = 20

Apples after day 2 = 80 - 20 = 60

Apples she ate on the 3rd day = 1/2\times60 = 30

So, remaining apples that she throws out = 60 - 30 = 30

So, when Ming threw out 30 apples, she bought 100

When she throws out 6 apples, then she bought = 100 x 6 30 = 20

So, Ming bought 20 apples.